OpenStudy (diyadiya):

If the roots of \(ax^2+bx+c=0\) are in the ratio m:n then A.\(mna^2=(m+n)c^2\) B.\(mnb^2=(m+n)ac\) C.\(mnb^2=(m+n)^2ac\) D.None of these

OpenStudy (diyadiya):

Let α and β be the roots \[\frac{α}{β}= \frac{m}{n}\]

OpenStudy (anonymous):

\[n \alpha = \beta m\]\[\alpha +\beta = \frac{-b}a\]\[\alpha \beta = \frac c a \]Use these.

OpenStudy (anonymous):

Substitute the values of the roots using the ratio relation, just do it. I don't wanna type it out all out, it's too algebraic to type.

OpenStudy (diyadiya):

Wait i didn't get that ..

OpenStudy (diyadiya):

what's the value of the roots here?

OpenStudy (callisto):

It's very messy :( Only things on the left are related to this question...

OpenStudy (callisto):

And for the second last step.. i should cross out the power of n, that is 2, instead of n, sorry :S

OpenStudy (diyadiya):

can you type the first equation ? :S

OpenStudy (callisto):

the first equation?

OpenStudy (diyadiya):

yeah i cant understand from the pic u attached

OpenStudy (anonymous):

lol, 'This sheet must not be taken away"

OpenStudy (callisto):

Sorry... Dunno if it makes better or worse

OpenStudy (diyadiya):

WOw Thank you Soo much!!!

OpenStudy (callisto):

Do you understand ?

OpenStudy (diyadiya):

Yup :)

OpenStudy (callisto):

Glad to help :D